The ratio of areas within the electron orbits for the first excited state to the ground state for the hydrogen atom is
1
Expert's answer
2018-03-09T10:52:07-0500
Radius of the n orbit rn ∝ n2.
For ground state n = 1 and for the first excited state n = 2.
Therefore
r2 / r1 = 22 / 1 = 4 / 1 = 4:1